a pottery manufacturer has an order to manufacture 5000 hanging vases to hold ½ pt. of water when full. The vases are designed so as to fit into the corner of a room. The faces of each vase are triangular in shape and intersect to form a pointed bottom. The area of the polygon cut out of the plane of the base by the lateral faces is 3sq.in. The height of the vase is 8 in. Compute the weight of pottery required if pottery weighs 130lb. per cu.ft.


1
Expert's answer
2021-11-04T15:55:11-0400
"V_ {out}= \dfrac{1}{3}Sh=\dfrac{1}{3}(3in^2)(8in)=8in^3"

Vase holds ½ pt. of water when full.

"V_{in}=1/2\ pt=14.4375in^3"

Since "V_ {out}=8in^3<14.4375in^3=V_{in}," then there is no solution.

Such vase is impossible.


Let us redefine the task:

a pottery manufacturer has an order to manufacture 5000 hanging vases to hold 1/4 pt. of water when full. The vases are designed so as to fit into the corner of a room. The faces of each vase are triangular in shape and intersect to form a pointed bottom. The area of the polygon cut out of the plane of the base by the lateral faces is 3sq.in. The height of the vase is 8 in. Compute the weight of pottery required if pottery weighs 130lb. per cu.ft.


"V_ {out}= \dfrac{1}{3}Sh=\dfrac{1}{3}(3in^2)(8in)=8in^3"

Vase holds 1/4 pt. of water when full.


"V_{in}=1/4\ pt=7.21875in^3"

Then


"V=V_{out}-V_{in}"

"=8in^3-7.21875in^3=0.78125in^3"

"1 lbs\/ft ^3=453.59237 g\/(1728in^3)"

"m_{vase}=130(\dfrac{453.59237 g}{1728in^3})(0.78125in^3)=26.6597g"

"M=5000(26.6597g)=133298.540g"

The total weight of the 5000 vases is 133 kg 299 g.

In this case we obtain the answer.